rubbing alcohol is typically diluted with water to 70% strength. if you need 3.5 ounces of 45% rubbing alcohol, how many ounces of 70% rubbing alcohol and how much water should you combine?
Answer by ptaylor(1894)


Let x=amount of 70% rubbing alcohol needed
Then 3.5-x=amount of water needed
Now we know that the amount of pure alcohol in the 70% solution (0.70x) plus the amount of pure alcohol in the water (0) has to equal the amount of pure alcohol in the final mixture (3.5*0.45). So our equation to solve is:
0.70x=3.5*0.45 or
0.70x=1.575 divide each side by 0.70
x=2.25 ounces-----------------------------amount of 70% alcohol needed
3.5-x=3.5-2.25=1.25 ounces---------------- amount of water needed
CK
2.25*0.70=3.5*0.45
1.575=1.575
Hope this helps---ptaylor